OpenStudy (mpj4):

How to get x = -3 and y = 5 using these equations? 3x-y+14=0 121+4xy-6x-28y^2+y^2+4x^2=0 -14x -72y+149-4xy+4y^2+x^2=0

OpenStudy (mpj4):

Ok so the problem is the circle passes through 1 point (-1,6) and tangent to two lines (2x+y+6=0) (x-2y+8=0) and I got values 3x-y+14=0, 121+4xy-6x-28y^2+y^2+4x^2 and -14x -72y+149-4xy+4y^2+x^2

OpenStudy (mpj4):

but I have no Idea how I'm supposed to work these out, I plugged it in a graphin thingy and I got the correct coordinates with them

OpenStudy (mpj4):

I'm trying to find the center btw

OpenStudy (mpj4):

the center I got is (-3,5)

OpenStudy (cwrw238):



OpenStudy (mpj4):

Here's the graphical representation of the equations I got, I just need to know how to solve for x and y https://www.desmos.com/calculator/cl7wozqmvg

OpenStudy (cwrw238):

the equation of the tangents go through A and B - then you can deduce the equations of the radii passing through the center O since they are perpendicular to the tangents. Solving these 2 equations s will give you the coordinates of the center

OpenStudy (cwrw238):

oh - you also need to know the coordinates of A and B ....

OpenStudy (cwrw238):

I'm not sure if you have enough info to do this question

OpenStudy (cwrw238):

sorry i gotta go right now...

OpenStudy (mpj4):

ah yeah, I got the equations actually 3x-y+14=0 121+4xy-6x-28y^2+y^2+4x^2=0 -14x -72y+149-4xy+4y^2+x^2=0 solving any 2 of these will get you x and y so says my professor, but I really don't know how.

OpenStudy (mpj4):

This is what those three equations should look like, but I really don't know how to get (-3,5)

ganeshie8 (ganeshie8):

heyy still stuck on this ?

OpenStudy (mpj4):

yes ;~;

ganeshie8 (ganeshie8):

Let the equation of required circle be \((x-h)^2 + (y-k)^2 = r^2\) since this circle passes through \((-1, 6)\), we have : \[(-1-h)^2 + (6-k)^2 = r^2 \tag{1}\]

ganeshie8 (ganeshie8):